# Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods

### v3

@@ 2.3 Discrete Fourier Transform (DFT)
We search for signals in the simulated data using the DFT version formulated by [19]. This method searches for the best pure sine model for the data. The notations for the DFT model are
@@
where $s_{\mathrm{DFT}}(t)$ is the pure sine signal and $p_{\mathrm{DFT}}(t)$ is the trend. The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively. Our DFT analysis may fail due to four application limitations.
The data sample is “too short” [22]
@@
The data contain “a trend” [26]
@@
The signal frequencies are “too close” [22].
@@
The signals are not “pure sines” [6].
@@
The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$ [26]. Then, the technique described by Ghaderpour et al. [15] is used to detect the pure sine signals. We apply the DFT to the detrended $y_{i}-p_{i}$ data. When searching for the second signal, we apply the DTF to the original model residuals $\epsilon_{i}=y_{i}-g_{\mathrm{y,DFT}}(t_{i})$.
@@ 3 Results
We apply DCM and DFT to seven different simulated data sets (Sections 3.1 - 3.7). As we proceed, these data sets become increasingly complex. DCM analysis succeeds for all data sets. DFT analysis fails for every data set.
@@ 3.1 Model 1
Our first simulation model is the one signal model
@@
Model 1 is a DCM model, where $K_{1}=1$, $K_{2}=1$, $K_{3}=0$, $B_{1,1}=$ $(A_{1}/2)\cos{(2\pi f_{1}t_{\mathrm{1,max,1}})}$ and $C_{1,1}=$ $(A_{1}/2)$ $\sin{(2\pi f_{1}t_{\mathrm{1,max,1}})}.$ We give the DCM analysis results for three samples having different $n$ and $\mathrm{SN}$} combinations (Table 1: Columns 2-4). For each sample, this table specifies the electronic data file and the electronic DCM control file. [^2] We publish all our data files and all DCM control files in https://zenodo.org/uploads/17018676 The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$. These values become more accurate when $n$ and $\mathrm{SN}$} increase.
A graphical presentation of DCM analysis results is shown for Model 1 simulated data sample, where $n=50$ and ${\mathrm{SN}}=50$ (Figures 1a-d). The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 1a: diamond). The DCM short search periodogram $z_{1}(f_{1})$ gives the best period at $P_{1}=1.822$ (Figure 1b: diamond). The continuous black line denoting the DCM model $g(t)$ fits perfectly to the black dots denoting the data $y_{i}$ (Figure 1c). The mean $p(t)=M_{0}=1.061\pm 0.036$ is correct (Figure 1c: dashed black line). The detrended model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots) and the pure sine signal $h_{1}(t)$ (red thick continuous line) are shown in Figure 1d. Note that the thick continuous red line stays under the thin continuous black line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -0.65 (blue dotted line). These residuals are stable and display no trends.
@@
For the simulated data of Model 1, the DCM analysis succeeds, but the DFT analysis fails.
@@ 3.2 Model 2
Our next one signal data simulation model is
@@
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$. We give the $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values in Table 2. As a DCM model, the orders of Model 2 are $K_{1}=1$, $K_{2}=1$ and $K_{3}=2$. The simulated data sample is “too short” because the period $P_{1}$ is $1.9\times\Delta T$ (Equation 27). The parabolic trend $p(t)$ is unknown (Equation 28). Again, we use the DCM and DFT time series analysis methods to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.63$ and $P_{\mathrm{max}}=3P_{1}=4.70$.
The DCM analysis results are given in Table 2. These results are not very accurate for the $n=100$ and ${\mathrm{SN}}=100$ combination, but they definitely improve for larger $n$ and $\mathrm{SN}$} values. For Model 2, the DCM analysis results are illustrated for the n=10000n=10~000 and ${\mathrm{SN}}=100$ combination (Figures 2a-d). The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 2a: diamond). The DCM short search gives the value $P_{1}=1.852$ (Figure 2b: diamond). The DCM model $g(t)$ is so good that it’s continuous black line is totally covered by the black dots representing the $y_{i}$ data (Figure 2c). Therefore, the colour of this $g(t)$ line has been changed from black to white. The results for the parabolic trend coefficients $M_{0}=1.079\pm 0.096$, $M_{1}=0.229\pm 0.026$ and $M_{2}=0.451\pm 0,064$ of the dashed black $p(t)$ line are correct. In Figure 1d, the white continuous line shows the detrended model g(t)p(t)g(t)-p(t). The black dots show the detrended data $y(t_{i})-p(t_{i})$ and the red thick continuous line shows the pure sine signal $h_{1}(t)$. Note that the red thick line is under the white thin line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -0.65. The colour of dotted line, which denotes this offset level, has been changed from blue to white. The distribution of these DCM model residuals is stable, as expected for a random normal distribution.
@@
Only DCM (not DFT) succeeds in the analysis of Model 2 simulated data.
@@ 3.3 Model 3
Our third data simulation model is the two signal model
@@
We give the $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$ and $M_{0}$ values in Table 3 (Column 1). The DCM orders of Model 3 are $K_{1}=2$, $K_{2}=1$ and $K_{3}=0$. The simulated data sample is not “too short” (Equation 27) because $P_{1}<P_{2}<\Delta T$. The $p(t)$ trend mean level $M_{0}$ is unknown (Equation 28). The two frequencies are “too close” because $\Delta f=f_{1}-f_{2}=f_{0}/2.72$ (Equation 29). We apply the DCM and DFT time series analyses to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
The DCM analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 3 (Columns 2-4). The results are surprisingly accurate even for the small $n=50$ sample having a low ${\mathrm{SN}}=10$. The DCM results for Model 3 are illustrated in Figures 1a-d ($n=50$ and ${\mathrm{SN}}=50$ combination). The DCM long search $z_{1}(f_{1})$ periodogram (red) and $z_{2}(f)$ periodogram (blue) minima are at the best periods $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 3a: diamonds). The DCM short search gives $P_{1}=0.160$ and $P_{2}=0.170$ (Figure 3b: diamonds). The black continuous DCM model $g(t)$ curve crosses trough the black dots of $y_{i}$ data (Figure 1c). The result $M_{0}=0.9985\pm 0.0032$ for the dashed black $p(t)$ trend line is correct. Our Figure 3d shows the detrended DCM model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the pure sine signal $h_{2}(t)$ (blue thin continuous line). The DCM residuals (blue dots) are offset to the level of -3.0 (blue dotted line). These residuals are small and their level is stable.
@@
The DCM analysis of Model 3 simulated data succeeds, but the DFT analysis does not.
@@ 3.4 Model 4
The next data simulation model is
@@
Model 4 is a DCM model having orders $K_{1}=2$, $K_{2}=1$ and $K_{3}=2$. Our DCM analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 4 (Columns 2-4). The DCM detects the correct $P_{1}$, $P_{2}$, $A_{1}$, $A_{2}$, $t_{\mathrm{max,1}}$, $t_{\mathrm{max,2}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination. Our Figures 4a-d illustrate the DCM analysis results for one sample of simulated data (Model 4: $n=50$ and ${\mathrm{SN}}=50$). The DCM long search best periods are at $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 4a: diamonds). The DCM short search values are $P_{1}=0.160$ and $P_{2}=0.170$ (Figure 4b: diamonds). The continuous DCM model $g(t)$ black line crosses through all black dots representing the $y_{i}$ data (Figure 4c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.9997\pm 0.0030$, $M_{1}=0.2565\pm 0.0066$ and $M_{2}=0.511\pm 0.011$ (Figure 4c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the pure sine signal $h_{2}(t)$ (blue thin continuous line) are shown in Figure 4d. The DCM residuals (blue dots) offset to the level of -3.0 show no trends and are extremely stable.
@@
Our Model 4 simulated data analysis succeeds for the DCM and fails for the DFT.
@@ 3.5 Model 5
The mathematical equation
@@
for our fifth model for simulated data is the same for Model 4 (Equation 3.4). However, this new Model 5 differs from the earlier Model 4 because we use totally different $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values (Table 5, Column 1). The simulated data sample is “too short” for both $P_{1}$ and $P_{2}$ periods (Equation 27). There is “a trend” $p(t)$ (Equation 28). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” (Equation 29). The three main reasons that can cause the failure of DFT analysis are present. We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.47$ and $P_{\mathrm{max}}=3P_{1}=4.20$.
Model 5 has DCM orders $K_{1}=2$, $K_{2}=1$ and $K_{3}=2$. We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 5 (Columns 2-4). The DCM fails to detect the correct $P_{1}$, $P_{2}$, …, $M_{1}$ and $M_{2}$ values for the the lowest n=10000n=10~000 and ${\mathrm{SN}}=1~000$ combination (Table 5, Column 2). This shows that the DCM can fail, just like any other time series analysis method, if the quality of data is too low. The DCM results for higher $n$ and $\mathrm{SN}$} combinations are correct (Table 5, Columns 3-4).
We show the DCM analysis results for one sample of Model 5 simulated data (Figures 5a-d: n=10000n=10~000 and ${\mathrm{SN}}=10~000$). The long and short DCM searches give $P_{1}=1.407$ and $P_{2}=1.917$ (Figure 5a: diamonds), and $P_{1}=1.393$ and $P_{2}=2.024$ (Figure 5b: Diamonds). The continuous line denoting the model $g(t)$ is totally covered by the black dots of $y_{i}$ data (Figure 5c). Therefore, we use white colour to highlight this DCM model $g(t)$ line. The detected polynomial trend $p(t)$ coefficients $M_{0}=0.81\pm 0.16$, $M_{1}=0.230\pm 0.013$ and $M_{2}=0.596\pm 0.084$ are correct (Figure 5c: dashed black line). We show the detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) in Figure 5d. The DCM residuals (blue dots) are offset to the level of -1.5. These blue dots appear black because there are 10 000 of them. For obvious reasons, the -1.5 level of these residuals is highlighted by a white dotted line. The DCM model residuals are stable and show no trends.
@@
The DCM analysis succeeds for Model 5 simulated data, but the DFT analysis fails.

### v4

@@ 2.4 Discrete Fourier Transform (DFT)
The DFT is one of the most prevalent frequency-domain parametric time series analysis methods for unevenly spaced time series. It searches for the best pure sine model for the data. The equivalent DCM model has the orders $K_{1}=1,K_{2}=1$ and $K_{3}=0$. Any time series analysis method must be remarkable if it performs better than the distinguished DFT. Therefore, we will compare how DCM and DFT perform in the analyses of simulated time series (Sections 3.1-3.7). We search for signals in these simulated time series by using the frequently applied DFT version formulated by [21] [^2] This paper had over 2600 citations in December 2025. Our notation for the DFT test statistic is $z_{\mathrm{PDF}}(f)$ [21, Equation 1]. The notations for the DFT model are
@@
where $s_{\mathrm{DFT}}(t)$ is the sum of pure sine signals and $p_{\mathrm{DFT}}(t)$ is the trend. The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively. Our DFT analyses of simulated time series may fail due to AL6-AL10.
The time series is “too short” [25, AL6]
@@
The time series is non-stationary because it contains “a trend” [33, AL7]
@@
Due to the leakage caused by the sample window (AL8), the signal frequencies are “too close” [25, AL9].
@@
The signals are not “pure sines” [5, AL10]
@@
Many parametric frequency-domain time series analysis methods, like the DFT, can be applied only to stationary data. Trends changing the time series mean or variance must be removed before applying these methods [33]. The DFT can detect only one period at the time. Such frequency-domain parametric time series analysis methods are hereafter called “one-dimensional”. Our DFT analysis of simulated data proceeds through two stages. First, the $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is removed from the simulated data $y_{i}=y(t_{i})$. Then, the iterative pre-whitening technique described by Ghaderpour et al. [17] is applied to search for the pure sine signals. We search for the first signal by applying the DFT to the detrended $y_{i}-p_{i}$ simulated data. In the second signal search, the DTF is applied to the original model residuals $\epsilon_{i}=(y_{i}-p_{i})-g_{\mathrm{y,DFT}}(t_{i})$. At any stage, this combination of detrending and iterative pre-whitening analysis may fail and corrupt the results obtained at the next stages. For example, all DFT analyses of simulated time series will fail already at the detrending stage (Sections 3.1-3.7). Our DCM has no corrupting separate stages because the trend and the signal(-s) are detected (modelled) simultaneously.
@@ 3 Results
We apply DCM and DFT to seven different simulated time series (Sections 3.1 - 3.7). As we proceed, these time series become increasingly complex. DCM analysis succeeds for all time series. DFT analysis fails for every time series.
@@ 3.1 Model 1
Our first time series simulation model is the one signal model
@@
Model 1 is a DCM model, where $K_{1}=1$, $K_{2}=1$, $K_{3}=0$, $B_{1,1}=$ $(A_{1}/2)\cos{(2\pi f_{1}t_{\mathrm{1,max,1}})}$ and $C_{1,1}=$ $(A_{1}/2)$ $\sin{(2\pi f_{1}t_{\mathrm{1,max,1}})}.$ We give the DCM analysis results for three samples having different $n$ and $\mathrm{SN}$} combinations (Table 1: Columns 2-4). For each sample, this table specifies the electronic data file and the electronic DCM control file. [^3] We publish all our data files and all DCM control files in https://zenodo.org/uploads/17018676 The detected $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$ and $M_{0}$ values are correct and accurate even for the combination $n=50$ and ${\mathrm{SN}}=50$. Regardless of $\Delta T<P_{1}$, these model parameter values become more accurate and converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase. This confirms the WD-effect.
A graphical presentation of DCM analysis results is shown for Model 1 simulated time series, where $n=50$ and ${\mathrm{SN}}=50$ (Figures 1a-d). The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 1a: diamond). The DCM short search periodogram $z_{1}(f_{1})$ gives the best period at $P_{1}=1.822$ (Figure 1b: diamond). The continuous black line denoting the DCM model $g(t)$ fits perfectly to the black dots denoting the data $y_{i}$ (Figure 1c). The mean $p(t)=M_{0}=1.061\pm 0.036$ is correct (Figure 1c: dashed black line). The detrended model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots) and the pure sine signal $h_{1}(t)$ (red thick continuous line) are shown in Figure 1d. Note that the thick continuous red line stays under the thin continuous black line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -0.65 (blue dotted line). These residuals are stable and display no trends.
@@
For the simulated time series of Model 1, the DCM analysis succeeds, but the DFT analysis fails.
@@ 3.2 Model 2
Our next one signal time series simulation model is
@@
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$. We give the $P_{1}$, $A_{1}$, $t_{\mathrm{1,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values in Table 2. As a DCM model, the orders of Model 2 are $K_{1}=1$, $K_{2}=1$ and $K_{3}=2$. The simulated time series is “too short” because the period $P_{1}$ is $1.9\times\Delta T$ (Equation 35). The parabolic trend $p(t)$ is unknown (Equation 36). Again, we use the DCM and DFT time series analysis methods to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.63$ and $P_{\mathrm{max}}=3P_{1}=4.70$.
The DCM analysis results are given in Table 2. These results are not very accurate for the $n=100$ and ${\mathrm{SN}}=100$ combination, but they definitely improve for larger $n$ and $\mathrm{SN}$} values. The WD-effect ensures that the detected values converge to the correct simulated model parameter values. The short sample window, $\Delta T<P_{1}$, does not mislead the DCM analysis.
For Model 2, the DCM analysis results are illustrated for the n=10000n=10~000 and ${\mathrm{SN}}=100$ combination (Figures 2a-d). The DCM long search $z_{1}(f_{1})$ periodogram minimum is at $P_{1}=1.843$ (Figure 2a: diamond). The DCM short search gives the value $P_{1}=1.852$ (Figure 2b: diamond). The DCM model $g(t)$ is so good that it’s continuous black line is totally covered by the black dots representing the $y_{i}$ data (Figure 2c). Therefore, the colour of this $g(t)$ line has been changed from black to white. The results for the parabolic trend coefficients $M_{0}=1.079\pm 0.096$, $M_{1}=0.229\pm 0.026$ and $M_{2}=0.451\pm 0,064$ of the dashed black $p(t)$ line are correct. In Figure 1d, the white continuous line shows the detrended model g(t)p(t)g(t)-p(t). The black dots show the detrended data $y(t_{i})-p(t_{i})$ and the red thick continuous line shows the pure sine signal $h_{1}(t)$. Note that the red thick line is under the white thin line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -0.65. The colour of dotted line, which denotes this offset level, has been changed from blue to white. The distribution of these DCM model residuals is stable, as expected for a random normal distribution.
@@
Only DCM (not DFT) succeeds in the analysis of Model 2 simulated time series.
@@ 3.3 Model 3
Our third time series simulation model is the two signal model
@@
We give the $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$ and $M_{0}$ values in Table 3 (Column 1). The DCM orders of Model 3 are $K_{1}=2$, $K_{2}=1$ and $K_{3}=0$. The simulated time series is not “too short” (Equation 35) because $P_{1}<P_{2}<\Delta T$. The $p(t)$ trend mean level $M_{0}$ is unknown (Equation 36). The two frequencies are “too close” because $\Delta f=f_{1}-f_{2}=f_{0}/2.72$ (Equation 37). We apply the DCM and DFT time series analyses to search for periods between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
The DCM analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 3 (Columns 2-4). The results are surprisingly accurate even for the small $n=50$ sample having a low ${\mathrm{SN}}=10$. Due to the WD-effect, the detected model parameter values converge to the correct simulated values when $n$ and $\mathrm{SN}$} increase.
The DCM results for Model 3 are illustrated in Figures 1a-d ($n=50$ and ${\mathrm{SN}}=50$ combination). The DCM long search $z_{1}(f_{1})$ periodogram (red) and $z_{2}(f)$ periodogram (blue) minima are at the best periods $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 3a: diamonds). The DCM short search gives $P_{1}=0.160$ and $P_{2}=0.170$ (Figure 3b: diamonds). The black continuous DCM model $g(t)$ curve crosses trough the black dots of $y_{i}$ data (Figure 1c). The result $M_{0}=0.9985\pm 0.0032$ for the dashed black $p(t)$ trend line is correct. Our Figure 3d shows the detrended DCM model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the pure sine signal $h_{2}(t)$ (blue thin continuous line). The DCM residuals (blue dots) are offset to the level of -3.0 (blue dotted line). These residuals are small and their level is stable.
@@
The DCM analysis of Model 3 simulated time series succeeds, but the DFT analysis does not.
@@ 3.4 Model 4
The next time series simulation model is
@@
Model 4 is a DCM model having orders $K_{1}=2$, $K_{2}=1$ and $K_{3}=2$. Our DCM analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 4 (Columns 2-4). The DCM detects the correct $P_{1}$, $P_{2}$, $A_{1}$, $A_{2}$, $t_{\mathrm{max,1}}$, $t_{\mathrm{max,2}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination. As the simulated data $n$ and $\mathrm{SN}$} values increase, the WD-effect ensures that all detected values converge to the correct simulated model parameter values.
Our Figures 4a-d illustrate the DCM analysis results for one sample of simulated time series (Model 4: $n=50$ and ${\mathrm{SN}}=50$). The DCM long search best periods are at $P_{1}=0.159$ and $P_{2}=0.168$ (Figure 4a: diamonds). The DCM short search values are $P_{1}=0.160$ and $P_{2}=0.170$ (Figure 4b: diamonds). The continuous DCM model $g(t)$ black line crosses through all black dots representing the $y_{i}$ data (Figure 4c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.9997\pm 0.0030$, $M_{1}=0.2565\pm 0.0066$ and $M_{2}=0.511\pm 0.011$ (Figure 4c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (black continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the pure sine signal $h_{2}(t)$ (blue thin continuous line) are shown in Figure 4d. The DCM residuals (blue dots) offset to the level of -3.0 show no trends and are extremely stable.
@@
Our Model 4 simulated time series analysis succeeds for the DCM and fails for the DFT.
@@ 3.5 Model 5
The mathematical time series equation
@@
for our fifth model is the same as for Model 4 (Equation 3.4). However, this new Model 5 differs from the earlier Model 4 because we use totally different $P_{1}$, $P_{2}$, $A_{1}$$A_{2}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{2,max,1}}$, $M_{0}$, $M_{1}$ and $M_{2}$ values (Table 5, Column 1). The simulated time series is “too short” for both $P_{1}$ and $P_{2}$ periods (Equation 35). There is “a trend” $p(t)$ (Equation 36). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” (Equation 37). The three main reasons that can cause the failure of DFT analysis are present. We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.47$ and $P_{\mathrm{max}}=3P_{1}=4.20$.
Model 5 has DCM orders $K_{1}=2$, $K_{2}=1$ and $K_{3}=2$. We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 5 (Columns 2-4). The DCM fails to detect the correct $P_{1}$, $P_{2}$, …, $M_{1}$ and $M_{2}$ values for the the lowest n=10000n=10~000 and ${\mathrm{SN}}=1~000$ combination (Table 5, Column 2). This shows that the DCM can fail, just like any other time series analysis method, if the quality of data is too low. Due to the WD-effect, the DCM results for higher $n$ and $\mathrm{SN}$} combinations are correct (Table 5, Columns 3-4).
We show the DCM analysis results for one sample of Model 5 simulated time series (Figures 5a-d: n=10000n=10~000 and ${\mathrm{SN}}=10~000$). The long and short DCM searches give $P_{1}=1.407$ and $P_{2}=1.917$ (Figure 5a: diamonds), and $P_{1}=1.393$ and $P_{2}=2.024$ (Figure 5b: Diamonds). The continuous line denoting the model $g(t)$ is totally covered by the black dots of $y_{i}$ data (Figure 5c). Therefore, we use white colour to highlight this DCM model $g(t)$ line. The detected polynomial trend $p(t)$ coefficients $M_{0}=0.81\pm 0.16$, $M_{1}=0.230\pm 0.013$ and $M_{2}=0.596\pm 0.084$ are correct (Figure 5c: dashed black line). We show the detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) in Figure 5d. The DCM residuals (blue dots) are offset to the level of -1.5. These blue dots appear black because there are 10 000 of them. For obvious reasons, the -1.5 level of these residuals is highlighted by a white dotted line. The DCM model residuals are stable and show no trends.
@@
The DCM analysis succeeds for Model 5 simulated time series, but the DFT analysis fails.